OpenStudy (anonymous):

Find the length of the legs of a 45°- 45°-90° triangle with a hypotenuse of 4.

OpenStudy (shinalcantara):

 note that you have 45-45-90 this only implies that the legs are having the same length. using pythagorean theorem you'll have: \[c^2 = a^2 + b^2\] and since a is equal to b then \[c^2 = a^2 + a^2\] \[c^2 = 2a^2\] c=4 substituting the value of c: \[4^2 = 2a^2\] \[a^2 = \frac{ 16 }{ 2 }\] \[a^2 = 8\] \[a = \sqrt{8}\] \[a = \sqrt{4\times2}\] \[a = 2\sqrt{2}\]

OpenStudy (anonymous):

Thank you so much for explaining it completely !